# Guided Bayesian Optimization: Data-Efficient Controller Tuning with Digital Twin

### v1

@@ IV-C1 Percentage overshoot
where $y_{\text{max}}$ is the maximum response in the transition period, $\Delta r:=r_{2}-r_{1}$ is the step height, $M_{\text{p}}:=y_{\text{max}}-y(T)$ is the peak overshoot, and $T$ is the final time. Percentage overshoot presents the degree of stability and damping in the system [47].
@@ IV-C4 Integral of time-weighted absolute error
such that $\sum_{i=1}^{i=4}w_{i}=1$ and $w_{i}>0$ where $w_{i}$ are the normalized weights of our selected individual performance metrics. Different weights can be used to attain different trade-offs between the performance metrics; in the experiments below, we measure each performance metric using $10$ controller gains evenly spanning the entire feasible set $\Theta$, then take the average of each metric as a normalization weight. We also adjust the weights based on the significance of each metric relative to the other. For example, we put more relevant gravity on minimizing overshoot related to performance and stability. The normalized weights $w_{1},w_{2},w_{3},w_{4}$ are selected as $0.44,0.22,0.22,0.11$, respectively, summing up to unit value.
The following Section will analyze our guided BO properties on $\mathrm{G}$.
@@ V Numerical Analysis
We numerically study the guided BO properties using the controller (15) and the identified plant in (14) comparing with the nominal controller tuning. Computations are performed by the Euler high-performance cloud-based cluster of ETH Zurich [50]. We use $8$ AMD EPYC 7742 CPUS cores with a clock speed of $2.25$ GHz nominal and $3.4$ GHz peak and $3072$ MB of memory.
The reference input $r$ of the closed-loop system equals the angular encoder speed $\omega$ in radian per second. We consider low and high reference encoder speed inputs equal to $r_{1}=245$ rad/s and $r_{2}=306$ rad/s corresponding to the nominal encoder speeds of $\mathrm{v}=1600$ and $\mathrm{v}=2000$, respectively. We model the output measurement noise by an additive Gaussian noise with zero mean and standard deviation $\sigma_{\epsilon}=0.03$. In the guided BO algorithm, we choose the RMSE threshold $\eta_{2}=3\sigma_{\epsilon}$ so that we accept DT plants with reasonably small RMSE considering the noise boundary according to (12b).
@@
Section System $\eta_{1}$ $\eta_{2}$ $\eta_{3}$ $N_{0}$ $\tilde{n}_{\text{EI}}$
V-A numerical 0$\infty$ 0.09-2.5 $\infty$ 1 5-40
V-B numerical 0$\infty$ 0.09 $\infty$ 1-10 5
V-C numerical $10^{-6}$-20 0.09 0.2 1 3
V-D numerical 3 0.09 0.2 1 3
VI-A rotary motor 3 0.09 0.2 1 3
VI-B linear motor 3 0.003 0.2 1 3
@@ V-A Digital Twin Fidelity
We run a Monte Carlo analysis of the guided BO algorithm using both DTs with $100$ batches of experiments. Each batch includes $50$ experiments on the real plant starting from an initial data set with a single data $N_{0}=1$. The controller gains of the initial data set are randomly selected inside the feasible set. So, the initial data sets of the batches are different from each other. Starting from the initial training data set, we set $\eta_{1}=0$ in (12a) and perform BO on the DT plant $\tilde{n}_{\text{EI}}$ times. Each time we activate DT, we continue BO iterations on DT up to a fixed number of iterations $\tilde{n}_{\text{EI}}$. Next, we deactivate the DT by setting $\eta_{1}=\inf$, and continue $N_{\mathrm{G}}$ iterations on the real plant until we activate the DT plant by setting $\eta_{1}=0$. This way, we activate DT periodically after every $N_{\mathrm{G}}$ iteration on the real plant $\mathrm{G}$. We set $\eta_{2}=0.09$ and $\eta_{2}=2.5$ when using high- and low-fidelity DT plants, respectively.
TABLE IV: Monte Carlo results with $100$ batches of $50$ experiments on the real plant, each starting from a different initial data set $\mathcal{D}$. Columns $3-5$ report the rounded average number of experiments over the real plant required to outperform the nominal optimality ratio $\phi=1.49$.
@@ V-C Digital Twin Activation Threshold
We determine the DT activation threshold $\eta_{1}$ in (12a). We fix the BO iterations on the real plant to $50$ experiments. At each threshold $\eta_{1}$, we repeat the guided BO algorithm $100$ times using different initial data sets. Fig. 4 shows that guided BO tunes the controller gains faster than BO when the DT is sufficiently activated, allowing the average optimality ratio of the measured performance to be low enough. More precisely, we observe in Fig. 4a that for guided BO, choosing a DT activation threshold in the range $\eta_{1}\in[2,8]$ (corresponding to DT activation around $12$ times on average in each batch) results in improved convergence of approximately $9$ experiments to the nominal performance compared to the BO method. We also observe that with large $\eta_{1}=20$, the guided BO approaches the BO’s performance indicated by dashed lines.
@@
Fig. 4b indicates the correlation between the total DT activations and the average optimality ratio measured on the real plant using various $\eta_{1}$ thresholds. For the DT activation threshold $\eta_{1}\geq 5$ in 12a, increasing $\eta_{1}$ decreases the number of times DT guides the optimizer and increases the average optimality ratio measured on the real plant. However, the cost estimated by DT is inaccurate, and over-activating the DT results in lower performance. That is why using small $\eta_{1}=1e-6$ results in the excessive activation of the DT, in which the average optimality ratio also increases to $2.2$. Therefore, choosing the DT activation threshold in the range $\eta_{1}\in[2,8]$ provides a better trade-off between the DT activation and the guided BO performance.
@@ V-D Performance Evaluation
A Monte Carlo study is performed with $100$ batches of experiments, each with fixed $35$ iterations on the real plant in (14). Table III specifies the guided BO algorithm parameter values. According to V-C, the parameter $\eta_{1}=3$ is chosen to provide us with an optimum trade-off between DT activation and the convergence speed of the algorithm. $\eta_{2}=0.09$ is chosen concerning three times the system noise boundary compared to the expected fidelity of the DT through RMSE error. We set the stopping threshold $\eta_{3}=0.2$ and $\tilde{n}_{\text{EI}}=3$ to avoid unnecessary optimization on the DT. The minimum number of initial data $N_{0}=1$ is chosen to provide the GP model with the least prior information when the guided BO is shown to improve the optimization further compared to BO as revealed in section V-B. The results are shown in Fig. 5. Guided BO with high-fidelity DT outperforms the nominal controller on average after approximately $2$ experiments, whereas BO requires approximately $12$ experiments.
@@
We compare the evaluated controller parameters of guided BO and BO in a batch of experiments starting from the same initial controller parameters $\Theta_{\text{init}}$. In Fig. 6, we visualize one batch of our Monte Carlo analysis consisting of $35$ control parameters proposed by each optimizer to measure the performance with the real plant. The true cost surface is presented with a colormap in Fig. 6a. This figure shows that the BO (magenta squares) requires exploring the feasible set. Fig. 6b depicts the cost estimated by DT after $34$ iterations on the real plant where the filled circles represent the measured parameters by guided BO on the real plant, and the hollow circles show the DT evaluations. We see that DT retrieves the true cost. Thus, the guided BO exploits the optimum region, avoiding the regions with higher costs.
@@
TABLE VI: Controller parameters tuned by BO and guided BO methods for the numerical system at iteration $n=21$.
Controller $\mathrm{K}_{\text{p}}$ $\mathrm{K}_{\text{i}}$
Ground truth optimum 0.54 1.16
Guided BO 0.53 1.16
BO 0.45 1.04
Fig. 8 shows the step response of the closed-loop system with the guided BO controller on the real plant, corresponding to batch $4$ and experiment $21$ in Fig. 5. Table VI specifies the corresponding tuned controller parameters in which the ground truth optimum controller gains are retrieved by dense grid search in the feasible set on the numerical system. While BO step responses have over-damped behavior, the guided BO outperforms the other data-driven and nominal approaches. The guided BO controller performance fits the ground truth performance with a slight overshoot, ITAE, and reasonably fast rise and settling time.
@@ VI-A Rotary Motor System
We want to tune the PI speed controller gains. We study the performance of our iterative tuning methods in the presence of unknown effects, such as friction and measurement noise. We focus on controlling the speed of the motor as it is relevant for its practical use. At the beginning of each experiment, we initialize the motor speed to reach a constant encoder angular speed equal to $r_{1}=245$ rad/sec. This allows us to remove the uncertainty of starting from an arbitrary initial condition on this hardware. We build the initial data set $\mathcal{D}$ with one gain pair ($N_{0}=1$) and the respective measured objective according to (23). With a sampling rate of $20$ Hz, we measure $\tilde{n}=100$ pairs of control commands $u$ and physical plant’s output $y$. In the guided BO algorithm, we use tfest function in MATLAB [51] with a second-order model to identify the DT of the nonlinear DC motor system based on refined instrumental variable method [39]. Table III summarizes the guided BO parameter parameter values.
We study the measured performance of $100$ batches $40$ experiments. Each batch starts with a different $\Theta_{\text{init}}$ randomly selected inside the feasible set $\Theta$ in (16). $\Theta_{\text{init}}$ used for each batch is identical for both guided BO and BO methods. We define the optimum cost as the minimum cost measured among all batches of experiments. The optimality ratio is between the minimum observed cost in each batch and the optimum cost.
Fig. 10 demonstrates that guided BO converges on average with fewer experiments on the real DC rotary motor system and has a smaller variance between repetitions. Let’s consider a threshold for the optimality ratio within $5\%$ of the optimal cost. BO requires, on average, $38$ experiments to converge to the optimality ratio of $1.05$, whereas guided BO requires only $11$ experiments to perform the same, a $71\%$ improvement. Indeed, the best among the $100$ batches of guided BO accomplishes this in $3$ experiments (compared to $7$ experiments required for BO, a $57\%$ improvement), and the worst in $30$ experiments (compared to $39$ experiments required by BO, a $23\%$ improvement). After $15$ experiments, we realized that all $100$ repetitions of guided BO attained optimality ratios better than $1.1$, compared to $55\%$ of the BO repetitions. In contrast, at this stage, the worst among the BO repetitions still had an optimality ratio of $2.85$. The dashed black line in Fig. 10 depicts the cost value given the first nominal PGM controller designed based on the identified linear model of the system as described in Section IV-A. Both data-driven controller tuning approaches outperform the nominal GPM method in terms of the cost value at the expense of the number of iterations needed.
@@
Fig. 11 depicts the histogram of all measured performance values. According to this histogram, BO requires performance measurements further away from the optimum, whereas the DT allows guided BO to test the controller gains at regions with smaller costs. Therefore, guided BO exploits more around the optimum cost, whereas BO explores further away from the optimum.
@@ VI-B Linear Servo Motor System
Our linear motor setup has a DM01 linear module from NTI AG company, displayed in Fig. 12. This module consists of a linear guide with an integrated LinMot P01 linear permanently actuated servo motor. We integrate a MS01-1/D-SSI non-contacting external position sensor with our module. This sensor has $5$ $\mu\text{m}$ absolute resolution and $0.005$ mm repeatability error. We build a unit feedback closed-loop system with a C1250-IP-XC-1S-C00 NTI AG axis controller and our external sensor to control the output position of the LinMot P01 linear motor with a sampling frequency of $1$ kHz. This industrial position controller is a PDT1 controller defined as
@@
where $\mathrm{K}_{\text{p}}$ and $\mathrm{K}_{\text{d}}$ are proportional and derivative gains, respectively. This controller commands the linear motor plant as a current signal measured in ampere. A saturation mechanism clips the control signal once it exceeds the range $[-7,7]$ ampere. Given the current and linear motor position measurements, our algorithm in MATLAB sets the reference output $r$ with certain industrial controller gains $\mathrm{K}_{\text{p}}$ and $\mathrm{K}_{\text{d}}$.
@@
We tune $\mathrm{K}_{\text{p}}$ and $\mathrm{K}_{\text{d}}$ gains of the position controller. To determine the RMSE threshold $\eta_{2}$ in (12b), we estimate the noise standard deviation by measuring the performance metrics in $5$ experiments with fixed identical controller gains. We normalize the standard deviation of each metric in $F=[\zeta,T_{\text{s}},T_{\text{r}},e_{\text{ITAE}}]$ according to their mean measured value to retrieve the inverse of the signal-to-noise ratio (1/SNR) equal to $[0.0001,0.004,0.002,0.001]$. Assuming weighted aggregation-based cost function in (23), we estimate the additive noise standard deviation as $\sigma_{\epsilon}=0.001$ mm. We choose the RMSE threshold $\eta_{2}=3\sigma_{\epsilon}=0.003$ to accept DT with sufficient fidelity considering the estimated noise level.
@@
We compare the performance of our proposed Guided BO with BO in $100$ batches of experiments, each with $25$ iterations on the physical plant. Each batch consists of initial data with $N_{0}$ and $25$ performance measurements on the real linear motor system, whereas Guided BO internally may estimate the cost function with its DT model. To build the initial data set $\mathcal{D}$ for the batches, we randomly select $100$ pair of gains $\theta=[\mathrm{K}_{\text{p}},\mathrm{K}_{\text{d}}]\in\Theta$. We use the same initial data set $\mathcal{D}$ for each batch in both BO and guided BO methods. In guided BO, each time that we estimate the performance with an updated DT model $\tilde{G}$, we remove previous $(\tilde{\theta},\tilde{\xi})$ data from the training data set $\mathcal{D}$. We perform an exhaustive grid search in $\Theta$ to calculate the ground truth optimum performance $J(\theta^{*})$ required to calculate the optimality ratios.
@@
Results on Fig. 13 prove that guided BO, on average, requires fewer experiments on the linear motor system and has a smaller variance between repetitions. Let’s put a threshold on the optimality ratio equal to $1.01$. BO requires, on average, $18$ experiments on the real plant, excluding the initial data set experiment, to converge to the optimality ratio of $1.01$.; in contrast, guided BO requires only $10$ experiments to perform the same, a $44\%$ improvement. Indeed, the best among the $100$ batches of guided BO accomplishes this in $2$ experiments (compared to $6$ experiments required for BO, a $66\%$ improvement), and the worst in $17$ experiments, compared to $24$ experiments required by BO, a $29\%$ improvement). After $10$ experiments, we realized that $86\%$ of guided BO iterations attained optimality ratios better than $1.01$, compared to only $14\%$ of the BO repetitions.
Figure 14: The histogram representing the percentage of experiments on the real system vs. the optimality ratio for $100$ batches of $25$ BO iterations. The guided BO exploits the regions with lower optimality ratios, whereas BO requires exploring regions with higher cost values.
Furthermore, the confidence interval of the guided BO shrinks faster than BO and both eventually shrink inside the noise boundary. Looking at the corresponding histogram of all measured performance on the actual system in Fig. 14 indicates that guided BO does not explore the controller parameters with high-cost values due to the guidance of its DT. However, BO alone requires further exploration, eventually leading to its slower tuning performance.
Table VII summarizes the optimum controller gains of both iterative controller tuning methods. This table validates the depicted step responses in Fig. 15 such that the derivative gain of the guided BO controller is higher than BO, causing smaller tracking error rate that results in a more minor overshoot, whereas the BO method has a larger proportional gain attempting to respond rapidly immediately after we change to a new reference signal. Furthermore, the optimum response has a maximum overshoot of $0.06\%$ lower than the required $1.12\%$ by the metric thread standard.
TABLE VII: Controller parameters tuned by BO and guided BO methods for the linear motor setup at iteration $i=13$ of a single batch of experiments in Fig. 13. The Optimum controller was retrieved from an exhaustive grid search.
@@
Finally, Fig. 15 visualizes the closed-loop response given the optimum controller gains of guided BO and BO methods after experiment $i=13$ of a batch of experiments. This figure indicates that the optimum step response tuned by guided BO converges faster than BO to the ground truth optimum response shown with the green line. We can rationalize the system’s behavior based on the aggregated performance metrics. Recalling our cost in (23), we notice that the area under the curve of guided BO response is less than BO. While the settling and rise times are similar, we realize that the guided BO reduces the overshoot, resulting in a lower aggregate cost value closer to the optimum response.
Figure 15: Closed-loop step response of the linear motor setup. Each figure corresponds to the optimum controller performance obtained after 1 to 4 iterations on the real system.

### v2

@@ IV-C1 Percentage overshoot
where $y_{\text{max}}$ is the maximum response in the transition period, $\Delta r:=r_{2}-r_{1}$ is the step height, $M_{\text{p}}:=y_{\text{max}}-y(T)$ is the peak overshoot, and $T$ is the final time. Percentage overshoot presents the stability degree and damping in the system [55].
@@ IV-C4 Integral of time-weighted absolute error
such that $\sum_{i=1}^{i=4}w_{i}=1$ and $w_{i}>0$ where $w_{i}$ are the normalized weights of our selected individual performance metrics. Different weights can be used to attain different trade-offs between the performance metrics; in the experiments below, we measure each performance metric using $5$ controller gains evenly spanning the entire feasible set $\Theta$, then take the average of each metric as a normalization weight.
We adjust the weights based on the relative significance of each metric. For instance, we put more relevant gravity on minimizing overshoot related to performance and stability. The normalized weights $w_{1},w_{2},w_{3},w_{4}$ are respectively selected as $0.44,0.22,0.22,0.11$, summing up to unit value.
The following Section will analyze our guided BO properties on the identified plant $\mathrm{G}$.
@@ V Numerical Analysis
We numerically study the guided BO properties using the controller (15) and plant $\mathrm{G}$ comparing with the nominal controller tuning. Computations are performed by the Euler high-performance cloud-based cluster of ETH Zurich [58]. We use $8$ AMD EPYC 7742 CPUS cores with a clock speed of $2.25$ GHz nominal and $3.4$ GHz peak and $3072$ MB of memory.
The reference input $r$ of the closed-loop system equals the angular encoder speed $\omega$ in radian per second. We consider low and high reference encoder speed inputs equal to $r_{1}=245$ rad/s and $r_{2}=306$ rad/s corresponding to the nominal encoder speeds of $\mathrm{v}=1600$ and $\mathrm{v}=2000$, respectively. We model the output measurement noise by an additive Gaussian noise with zero mean and standard deviation $\sigma_{\epsilon}=0.03$.
@@
Section System $\eta_{1}$ $\eta_{2}$ $\eta_{3}$ $N_{0}$ $\tilde{n}_{\text{EI}}$ $n_{\text{max}}$
V-A numerical $\infty 0$.09-2.5 $\infty$ 1 5-40 50
V-B numerical $\infty 0$.09 $\infty$ 1-10 5 50
V-C numerical $10^{-6}$-20 0.09 0.2 1 3 50
V-D numerical 3 0.09 0.2 1 3 35
VI-A rotary motor 3 0.09 0.2 1 3 40
VI-B linear motor 3 0.003 0.2 1 3 25
@@ V-A Digital Twin Fidelity
We run a Monte Carlo analysis of the guided BO algorithm using both DTs with $100$ batches of experiments. Each batch includes $n_{\text{max}}=50$ experiments on the real plant starting from an initial data set with a single data $N_{0}=1$. The controller gains of the initial data set are randomly selected inside the feasible set. So, the initial data sets of the batches are different from each other. Starting from the initial training data set, we set $\eta_{1}=0$ in (12a) and perform BO on the DT plant $\tilde{n}_{\text{EI}}$ times. Each time we activate DT, we continue BO iterations on DT up to a fixed number of iterations $\tilde{n}_{\text{EI}}$. Next, we deactivate the DT by setting $\eta_{1}=\inf$, and continue $N_{\mathrm{G}}$ iterations on the real plant until we activate the DT plant by setting $\eta_{1}=0$. This way, we activate DT periodically after every $N_{\mathrm{G}}$ iteration on the real plant $\mathrm{G}$. Here we set $\eta_{2}=0.09$ and $\eta_{2}=2.5$ when using high- and low-fidelity DT plants, respectively.
TABLE IV: Monte Carlo results with $100$ batches of $50$ experiments on the real plant, each starting from a different initial data set $\mathcal{D}$. Columns 353-5 report the rounded average number of experiments over the real plant required to outperform the nominal optimality ratio $\phi=1.49$.
@@ V-C Digital Twin Activation Threshold
We determine the DT activation threshold $\eta_{1}$ in (12a). At each threshold $\eta_{1}$, we repeat the guided BO algorithm $100$ times using different initial data sets. Fig. 4 shows that guided BO tunes the controller gains faster than BO when the DT is sufficiently activated, allowing the average optimality ratio of the measured performance to be low enough. More precisely, we observe in Fig. 4a that for guided BO, choosing a DT activation threshold in the range $\eta_{1}\in[2,8]$ (corresponding to DT activation around $12$ times on average in each batch) results in improved convergence of approximately $9$ experiments to the nominal performance compared to the BO method. We also observe that with large $\eta_{1}=20$, the guided BO approaches the BO’s performance indicated by dashed lines.
@@
Fig. 4 indicates the correlation between the total DT activations and the average optimality ratio measured on the real plant using various $\eta_{1}$ thresholds. For the DT activation threshold $\eta_{1}\geq 5$ in 12a, increasing $\eta_{1}$ decreases the number of times DT guides the optimizer and increases the average optimality ratio measured on the real plant. However, the cost estimated by DT is inaccurate, and over-activating the DT results in lower performance. That is why using small $\eta_{1}=1e-6$ results in the excessive activation of the DT, in which the average optimality ratio also increases to $2.2$. Therefore, choosing the DT activation threshold in the range $\eta_{1}\in[2,8]$ provides a better trade-off between the DT activation and the guided BO performance.
@@ V-D Performance Evaluation
A Monte Carlo study is performed with $100$ batches of experiments. Table III specifies the guided BO algorithm parameter values. According to V-C, the parameter $\eta_{1}=3$ is chosen to provide us with an optimum trade-off between DT activation and the convergence speed of the algorithm. $\eta_{2}=0.09$ is chosen concerning three times the system noise boundary compared to the expected fidelity of the DT through RMSE error. We set the stopping threshold $\eta_{3}=0.2$ and $\tilde{n}_{\text{EI}}=3$ to avoid unnecessary optimization on the DT. The minimum number of initial data $N_{0}=1$ is chosen to provide the GP model with the least prior information when the guided BO is shown to improve the optimization further compared to BO as revealed in section V-B. The results are shown in Fig. 5. Guided BO with high-fidelity DT outperforms the nominal controller on average after approximately $2$ experiments, whereas BO requires approximately $12$ experiments.
@@
We compare the evaluated controller parameters of guided BO and BO in a batch of experiments starting from the same initial controller parameters $\Theta_{\text{init}}$. In Fig. 6, we visualize one batch of our Monte Carlo analysis consisting of $35$ control parameters proposed by each optimizer to measure the performance with the real plant. The true cost surface is presented with a colormap in Fig. 6 which is a non-convex function of the controller parameters. This figure shows that the BO (magenta squares) requires exploring the feasible set. Fig. 6 (bottom panel) depicts the cost estimated by DT after $34$ iterations on the real plant where the filled circles represent the measured parameters by guided BO on the real plant, and the hollow circles show the DT evaluations. We see that DT retrieves the true cost. Thus, the guided BO exploits the optimum region, avoiding the regions with higher costs.
@@
TABLE VI: Comparison of optimum cost and controller parameters tuned by ground truth grid search, nominal, BO and guided BO methods after $n=21$ iterations on $\mathrm{G}$.
Controller $\mathrm{K}_{\text{p}}$ $\mathrm{K}_{\text{i}}$ $\hat{J}([\mathrm{K}_{\text{p}},\mathrm{K}_{\text{i}}])$
Ground truth grid search 0.54 1.16 0.79
Nominal($\Phi_{\text{m}}=60{\textdegree}$) 0.85 1.07 1.05
BO 0.45 1.18 0.84
Guided BO 0.53 1.16 0.79
Table VI specifies the tuned controller parameters and corrosponding cost values. The nominal controller is the PGM controller introduced in subsection IV-B with phase margin of $60{\textdegree}$. The ground truth optimum controller gains are retrieved by dense grid search in the feasible set on the numerical system. According to this table, the tuned controller by our guided BO method has lower cost than BO and the nominal controller, and its tuned controller parameters are almost equal to the ground truth optimum gains.
Fig. 8 shows the step response of the closed-loop system with the guided BO controller on the real plant, corresponding to batch $4$ and experiment $21$ in Fig. 5. While BO step responses have over-damped behavior, the guided BO outperforms the other data-driven and nominal approaches. However, the nominal has larger overshoot resulting from its large proportional gain. The guided BO controller performance fits the ground truth performance with a slight overshoot, ITAE, and reasonably fast rise and settling time.
@@ VI-A Rotary Motor System
We want to tune the PI speed controller gains. We study the performance of our iterative tuning methods in the presence of unknown effects, such as friction and measurement noise. We focus on controlling the speed of the motor as it is relevant for its practical use. At the beginning of each experiment, we initialize the motor speed to reach a constant encoder angular speed equal to $r_{1}=245$ rad/sec. This allows us to remove the uncertainty of starting from an arbitrary initial condition on this hardware. We build the initial data set $\mathcal{D}$ with one gain pair ($N_{0}=1$) and the respective measured objective according to (23). With a sampling rate of $20$ Hz, we measure $\tilde{n}=100$ pairs of control commands $u$ and physical plant’s output $y$. In the guided BO algorithm, we use tfest function in MATLAB [59] with a second-order model to identify the DT of the nonlinear DC motor system based on refined instrumental variable method [47]. By measuring five times the performance metrics in (23) with a randomnly selected $\theta\in\Theta$, we estimate the additive noise variance $\sigma_{\epsilon}=0.03$. Table III summarizes the guided BO parameter parameter values.
We study the measured performance of $100$ experiment batches. Each batch starts with a different $\Theta_{\text{init}}$ randomly selected inside the feasible set $\Theta$ in (16), which is identical for both guided BO and BO methods. We define the optimum cost as the minimum cost measured among all batches of experiments. The optimality ratio is between the minimum observed cost in each batch and the optimum cost.
@@
Fig. 10 demonstrates that guided BO converges on average with fewer experiments on the real DC rotary motor system and has a smaller variance between repetitions. Let’s consider a threshold for the optimality ratio within $5\%$ of the optimal cost. BO requires, on average, $38$ experiments to converge to the optimality ratio of $1.05$, whereas guided BO requires only $11$ experiments to perform the same, a $71\%$ improvement. Indeed, the best among the $100$ batches of guided BO accomplishes this in $3$ experiments (compared to $7$ experiments required for BO, a $57\%$ improvement), and the worst in $30$ experiments (compared to $39$ experiments required by BO, a $23\%$ improvement). After $15$ experiments, all $100$ repetitions of guided BO attain optimality ratios better than $1.1$, compared to $55\%$ of the BO repetitions. In contrast, at this stage, the worst among the BO repetitions still has an optimality ratio of $2.85$. The dashed black line in Fig. 10 depicts the cost value given the first nominal PGM controller designed based on the identified linear model of the system as described in Section IV-A. Both data-driven controller tuning approaches outperform the nominal GPM method in terms of the cost value at the expense of the number of iterations needed.
From the histogram of all measured performance values on Fig. 11, it can be seen that BO on the actual system requires performance measurements further away from the optimum, whereas the DT carries out the exploration allowing guided BO to test the controller gains on the real system at regions with smaller costs. Therefore, guided BO exploits more around the optimum cost, whereas BO explores further away from the optimum.
@@ VI-B Linear Servo Motor System
Our linear motor setup has a DM01 linear module from NTI AG company, displayed in Fig. 12. This module consists of a linear guide with an integrated LinMot P01 linear permanently actuated servo motor. We integrate a MS01-1/D-SSI non-contacting external position sensor with our module. This sensor has 55 $\mu\text{m}$ absolute resolution and $0.005$ mm repeatability error. We build a unit feedback closed-loop system with a C1250-IP-XC-1S-C00 NTI AG axis controller and our external sensor to control the output position of the LinMot P01 linear motor with a sampling frequency of $1$ kHz. This industrial position controller is a PDT1 controller implemented in C and defined as
@@
where $\mathrm{K}_{\text{p}}$ and $\mathrm{K}_{\text{d}}$ are proportional and derivative gains, respectively. This controller commands the linear motor plant as a current signal. A saturation mechanism clips the control signal once it exceeds the range $7\mathrm{A},7\mathrm{A}]$. We use OPC Unified Architecture in MATLAB to communicate with the hardware for the input and output measurements and controller parameter tuning. Given the current and linear motor position measurements, our algorithm sets the reference output $r$ with certain industrial controller gains $\mathrm{K}_{\text{p}}$ and $\mathrm{K}_{\text{d}}$.
@@
We tune $\mathrm{K}_{\text{p}}$ and $\mathrm{K}_{\text{d}}$ gains of the position controller. To determine the RMSE threshold $\eta_{2}$ in (12b), we estimate the noise standard deviation by measuring the performance metrics in $5$ experiments with fixed identical controller gains. We normalize the standard deviation of each metric in $\mathrm{f}=[\zeta,T_{\text{s}},T_{\text{r}},e_{\text{ITAE}}]$ according to their mean measured value to retrieve the inverse of the signal-to-noise ratio (1/SNR) equal to $[0.0001,0.004,0.002,0.001]$. Assuming weighted aggregation-based cost function in (23), we estimate the additive noise standard deviation as $\sigma_{\epsilon}=0.001$. We choose the RMSE threshold $\eta_{2}=3\sigma_{\epsilon}=0.003$ to accept DT with sufficient fidelity considering the estimated noise level.
@@
[…]
@@
Results on Fig. 13 prove that guided BO, on average, requires fewer experiments on the linear motor system and has a smaller variance between repetitions. Let’s put a threshold on the optimality ratio equal to $1.01$. BO requires, on average, $18$ experiments on the real plant, excluding the initial data set experiment, to converge to the optimality ratio of $1.01$.; in contrast, guided BO requires only $10$ experiments to perform the same, a $44\%$ improvement. Indeed, the best among the $100$ batches of guided BO accomplishes this in $2$ experiments (compared to $6$ experiments required for BO, a $66\%$ improvement), and the worst in $17$ experiments, compared to $24$ experiments required by BO, a $29\%$ improvement). After $10$ experiments, $86\%$ of guided BO iterations attain optimality ratios better than $1.01$, compared to only $14\%$ of the BO repetitions.
Figure 14: Percentage of experiments on the real system vs. the optimality ratio for $100$ batches in a fixed number ($25$) BO iterations.
Furthermore, the confidence interval of the guided BO shrinks faster than BO and both eventually shrink inside the noise boundary. Looking at the corresponding histogram of all measured performance on the actual system in Fig. 14, guided BO does not explore the controller parameters with high-cost values due to the guidance of its DT. However, BO alone requires further exploration, eventually leading to its slower tuning performance.
We summarize the optimum controller gains of guided BO, standard BO, and the ground truth in Table VII. The guided BO controller’s $\mathrm{K}_{\text{d}}$ is higher than the one obtained with BO, in line with the step responses shown in Fig. 15, causing smaller tracking error rate and a minor overshoot. The standard BO method has a larger $\mathrm{K}_{\text{p}}$ attempting to respond rapidly immediately after we change to a new reference signal. The optimum response has a maximum overshoot of $0.06\%$, lower than the required $1.12\%$ by the metric thread standard.
TABLE VII: Controller parameters for the linear motor setup at iteration $i=13$ of a single batch of experiments in Fig. 13.
@@
Figure 15 visualizes the closed-loop response given the optimum controller gains of guided BO and BO methods after experiment $i=13$ of a batch of experiments. This figure indicates that the optimum step response tuned by guided BO converges faster than BO to the ground truth optimum response shown with the green line. We can rationalize the system’s behavior based on the aggregated performance metrics. Recalling our cost in (23), we notice that the area under the curve of guided BO response is less than BO. While the settling and rise times are similar, we realize that the guided BO reduces the overshoot, resulting in a lower aggregate cost value closer to the optimum response.
Figure 15: Optimum controller performance at different iterations on the real system (linear motor closed-loop step response).